OpenStudy (anonymous):

x + 3 6 ----- = -- x + 2 5 solve for the varible x

OpenStudy (anonymous):

(x+3)5 = 6(x+2) 5x+15 = 6x+12 15-12 = 6x-5x 3 = x x = 3